A ball whose diameter is 5 cm is placed 80 cm in front of a convex spherical mirror. If the focal length of the mirror is -40 cm, what is the diameter of the ball as viewed in the mirror?
Answer #1

mirror equation

1/f = 1/u+1/v

1/v = 1/f-1/u

v = uf/u-f

v = 2.5*-40/42.5 = -2.35 cm

diameter = -4.7 cm